OpenStudy (anonymous):

find the domain y=f(x)=x^2-6x+6

hartnn (hartnn):

domain means the values which 'x' can take in that equation.

hartnn (hartnn):

do u see any values that 'x' can't take ?

OpenStudy (anonymous):

i dont get it......??

hartnn (hartnn):

(1)there is no denominator, if there was, then 'x' could not take values for which denominator=0

OpenStudy (anonymous):

so ther is no soln for this??

hartnn (hartnn):

(2) there is no \(\sqrt.\) sign, if there was , then 'x' cannot take values for which the expression under \(\sqrt{...}\)becomes negative.

hartnn (hartnn):

here, none of the 2 cases (1) or (2) arise. so 'x' can take all real values.

hartnn (hartnn):

so, domain is ALL real numbers R

hartnn (hartnn):

understood ?

OpenStudy (anonymous):

sorry i quite didnt get it

hartnn (hartnn):

which par ?

OpenStudy (anonymous):

x belongs to R then can we supoose any no we want in the place of x??

hartnn (hartnn):

yes, x can take any real value. u have options/choices ?

OpenStudy (anonymous):

how wud u solve this prob?

hartnn (hartnn):

example : f(x) =1/ x x cannot take value =0 f(x) = 1/(x-3) x cannot take value = 3 f(x) = sqrt{x-4} x cannot take value less than 4

OpenStudy (anonymous):

can u be mo` specific...?i cant understand

hartnn (hartnn):

i just gave u specific examples. and there's nothing to solve, you can say 'x' can take all real values...

OpenStudy (anonymous):

y did u add -3 in da denominator?

Parth (parthkohli):

By the way, all polynomials have the domain \((-\infty,\infty)\) a.k.a \(\mathbb{R}\).

OpenStudy (anonymous):

can u do the whole process??

OpenStudy (anonymous):

Domain means whenever the function exists. To figure this out, just find when the function does not exist, meaning when you're dividing by 0. In this case, you're never dividing by 0, so the function ALWAYS exists, so the domain is ]-inf,inf[, also know as "R" for all real numbers. If say f(x) = 1/(x-4), then the domain is ]-inf,-4[U]-4,inf[, because the function exists everywhere except at x=-4.